OpenStudy (anonymous):

The sides of an isosceles triangle have lengths 3x-1, -2x+29. The base has length 4x+5. What is the length of the base?

OpenStudy (anonymous):

help please!

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

the congruent legs are 3x-1 and -2x+29

jimthompson5910 (jim_thompson5910):

so 3x-1 = -2x+29 solve this for x

jimthompson5910 (jim_thompson5910):

once you have x, use it to find 4x+5

OpenStudy (anonymous):

That's how i thought you solve it, but each time i try to solve for x i get 3x=28 but if you divide28 by 3 you get a decimal, what am i doing wrong?

jimthompson5910 (jim_thompson5910):

3x-1 = -2x+29 3x-1+2x = -2x+29 +2x 3x+2x -1 = 29 3x+2x -1+1 = 29+1 3x+2x = 29+1 I'm sure you see where you went wrong at this point

OpenStudy (anonymous):

i see where i went wrong i forgot to write the x on -2 so the answer is 29(:

jimthompson5910 (jim_thompson5910):

yes because x = 6 base = 4x+5 base = 4*6 + 5 base = 24+5 base = 29 units

OpenStudy (anonymous):

thanks for helping!

jimthompson5910 (jim_thompson5910):

np